OpenStudy (yanasidlinskiy):

Please Help!!:D Thanx in advance:D Find the solutions of the equation. 1/2x^2 - x + 5 = 0

OpenStudy (anonymous):

x= 1+- sqrt(5)

OpenStudy (yanasidlinskiy):

That is NOT in any of my answer choice.

OpenStudy (yanasidlinskiy):

Ok. Can you show me how?

OpenStudy (yanasidlinskiy):

Is that the final answer? \(\huge{\uparrow}\)

OpenStudy (yanasidlinskiy):

Well that's not how its set up on my homework. I'm guessing that \(\huge{\uparrow}\) would be correct. It's set up like this: \[1/2x^2-x+5=0\]

OpenStudy (yanasidlinskiy):

It's ok. At least you gave it a try:D

OpenStudy (anonymous):

do you already have the answer or can i help?

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

x=-1-3,3-1

OpenStudy (aaronq):

Is it \(\dfrac{1}{2x^2} - x + 5 = 0 \) or \(\dfrac{x^2}{2} - x + 5 = 0 \) ?

undeadknight26 (undeadknight26):

yana banana no figure it out yet?

OpenStudy (yanasidlinskiy):

Hahha, sadly no.

undeadknight26 (undeadknight26):

I too lazy. I think i forgot how to do this also.

OpenStudy (yanasidlinskiy):

Lol, nvm. I actually got it right. I thought I was wrong:D Thank God!!! I went crazy all over this hoping to get it right. But anyways, thankx to all:D